OpenStudy (anonymous):

use fraction bars to find the quotient 2/9divde1/6 please tell me method

OpenStudy (anonymous):

\[(2/9)/(1/6)=(9*1)/(2*6)=9/12=(3*3)/(3*4)=3/4=0.75\]

OpenStudy (anonymous):

If you have to divide fractions like this: 1/2 divided by 1/3, it's equivalent to having to multiply 1/2 by 3/1. In general form, it's like this a/b divided by c/d = a/b multiplied by d/c. You just flip the fraction you're dividing by.